Waylander Coatings Company purchased waterproofing equipment on January 6 for $320,000. The equipment was expected to have a useful life of four years, or 20,000 operating hours, and a residual value of $35,000. The equipment was used for 7,200 hours during Year 1, 6,400 hours in Year 2, 4,400 hours in Year 3, and 2,000 hours in Year 4.

1. Determine the amount of depreciation expense for the years ended December 31, Year 1, Year 2, Year 3, and Year 4, by (a) the straight-line method, (b) the units-of-activity method, and (c) the double-declining-balance method. Also determine the total depreciation expense for the four years by each method. Note: FOR DECLINING BALANCE ONLY, round the multiplier to four decimal places. Then round the answer for each year to the nearest whole dollar. Depreciation Expense Year Year 1 Year 2 Year 3 Year 4 Total 2. What method yields the highest depreciation expense for Year 1? Straight-Line Method Units-of-Activity Method Double-Declining-Balance Methoc 3. What method yields the most depreciation over the four-year life of the equipment?
Answer #1

1 Ans:

Depreciation Expense

Years

Straight Line Method

Units of Output method

Double declining Balance Method

Year 1

$71,250

$102,600

$160,000

Year 2

$71,250

$91,200

$80,000

Year 3

$71,250

$62,700

$40,000

Year 4

$71,250

$28,500

$5000

Totals

$285,000

$285,000

$285,000

2) Ans:

Double declining balance methods yields the highest depreciation expense for Year 1 is $160,000

3) Ans:

All the methods Yield same amount of depreciation over the four year period equal to $285,000

Calculations:

Straight line method=$320,000-$35,000/4=$71,250

Units of Output method=$285,000*7200/20,000=$102,600: $285,000*6400/20,000=$91,200: $285,000*4400/20,000=$62,700: $285,000*2000/20,000=$28,500:

Double declining Balance Method=$320,000/4=80,000=80,000/320,000=0.25=.25*2=.50=50% =Double declining balance depreciation rate ,,,subject to the Total value of $285,000